OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 2x + y = 6 y = 3x + 4 What is the resulting equation?

OpenStudy (anonymous):

Are you looking for the final answer?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Then x=2/5 and y=26/5

OpenStudy (anonymous):

so y + 2x + y = 6?

OpenStudy (anonymous):

Well take the equation for y (y=3x+4) and plug it in to the top equation: 2x+(3x+4)=6

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

No problem

HanAkoSolo (jamierox4ev3r):

@MawellMcNulty make sure you show your work before giving out the answer. Other than that, good work!

OpenStudy (anonymous):

Oh my mistake I apologize